OpenStudy (anonymous):

how do i do this 0=2x^2+4x+3 need step by step please

OpenStudy (jhannybean):

An easier method to solving this would be completing the square.

OpenStudy (anonymous):

yah i did that, im just trying to find the x-intercept

OpenStudy (anonymous):

i know that y intercept is 3

OpenStudy (jhannybean):

So in solving this, you have \[2x^2+4x+3=0\]\[2(x^2+2x)=-3\]\[2(x^2+2x+\color{red}{1})=-3\color{red}{+2}\]\[2(x+1)^2=-1\]

OpenStudy (jhannybean):

In solving this, I first factored out a 2.

OpenStudy (jhannybean):

Because Im adding the same value to both sides of the equation, I have to keep in mind that im multiplying the left side of my equation by 2, so essentially i'm adding 2 to the right side. does that make sense?

OpenStudy (anonymous):

i got 2(x+1)^2+1

OpenStudy (jhannybean):

That is correct.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

alright thanks

OpenStudy (jhannybean):

And now solving for x

OpenStudy (jhannybean):

\[2(x+1)^2=-1\]\[(x+1)^2=\frac{-1}{2}\]\[x+1=\frac{\sqrt{-1}}{\sqrt{2}}\]\[x=~?\]

OpenStudy (jhannybean):

Recalling that \(i=\sqrt{-1}\)